OpenStudy (anonymous):

Rationalize the denominator 9/(7-3sqrt3)-(6-4sqrt3)

OpenStudy (dumbcow):

first simplify by combining like terms --> 9/(1+sqrt(3))

OpenStudy (anonymous):

\[\frac{9}{7-3\sqrt{3}-6+4\sqrt{3}}=\frac{9}{1+\sqrt{3}} * \frac{1-\sqrt{3}}{1-\sqrt{3}}\]

OpenStudy (anonymous):

multiply it out, the denominator will be rationalized. 1-sqrt(3) is the conjugate of 1+sqrt(3)